1. A ball is tossed straight up from a height of 17 meters with an initial speed of 7.2 m/s. What is its height above ground in meters after 0.869 seconds? Use g = 9.80 m/s2 . Assume 3 significant figures in your answer.   

2. Suppose you release an object from rest from a height of 7.08 meters. How long does it take the object to hit the ground if g = 9.80 m/s2? Assume 3 significant digits in your answer.
Answer #1

1.

H = h + ut - 0.5 gt^2

H = 17 + 7.2* 0.869 - 4.9* 869^2

H = 19.556 m

=======

2.

using 2nd equation of motion

7.08 = 0.5* 9.8* t^2

t= 1.202 s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck